# On the use of quantality in nuclei and many-body systems

### v1

@@ I Introduction
Quantality, a powerful dimensionless ratio first introduced by de Boer [1], was used by Mottelson to indicate when a system behaves like a quantum liquid (QL) rather than a crystal. It is larger for quantum liquid states than for crystal ones [2]. The former have constituents with delocalized wave-functions, leading to a homogeneous density, whereas the latter have localized constituents at given nodes. Superfluid Helium systems, nuclei, and atomic electrons are of QL type [2, 3, 4]. More precisely, nuclei and atomic electrons are found to be the two most QL-like systems, having larger values of the quantality than superfluid Helium systems.
@@
Despite its powerfulness, the role of quantality has been very scarcely studied in nuclear physics [2, 3, 13, 4]. In this work, we further explore how quantality can be used to understand the main differences between atomic electrons and nucleons in their quantum liquid phases. It is then used to identify the control parameters for the cluster phase in nuclei, namely the role of the number of nucleons and the density. The relation to the low-density transition towards clusterization is analyzed. The use of quantality, as a dimensionless ratio, to understand its relation with the interaction coupling constants leads to a fundamental relation, allowing to study the magnitude of the spin-orbit effect in finite nuclei. The relative magnitude of the binding energy to the shell-structure one is also shown to depend on the quantality.
The present work aims to analyze the behavior of many-body systems with simple but powerful indicators, by progressively refining their quantum description. First, the role of the Zero Point Energy (ZPE) is discussed. Then, the inclusion of the dispersion of the wave-function of the constituents, from the solution of a e.g. Schrödinger equation is considered. As a third step, the relation with the spin-orbit effect is given, based on the non-relativistic reduction of the Dirac equation. It will be shown how quantality plays a pivotal role during all these three stages. In section II, the emergence of a fundamental lengthscale, the limit radius, is studied in relation to the ZPE and the quantality. The relation between the quantality and the localization parameter is discussed in section III, showing how the appearance of clusters in the low-density regime is recovered. Section IV relates the quantality to the dimensionless magnitude of the interaction, showing how the spin-orbit effect in finite systems can be deduced. Finally, a systematic study of possible dimensionless quantities is provided in Section V, generalizing the concept of quantality.
@@ II The quantality, the zero point energy, and the limit radius
Let us consider a many-body system composed of constituents of mass m and interacting with an interaction of typical magnitude V0. The inter-constituent distance at equilibrium is r0. These 3 quantities characterize the main behavior of the system as they correspond to the input of its equation of motion. Instead of solving it as exactly as possible, a complementary way to predict the qualitative behavior of the system is to use indicators, incorporating step by step quantum mechanical effects. As a first step, the quantality is defined as the ratio of the ZPE T0 to the magnitude of the potential V0:
Quantality was used by Mottelson to characterize QL vs crystal behavior of systems [2]. The typical value calculated with Eq. (1) is 10-2,-3 in the case of crystals like atoms and molecules, and 0.1-1 (i.e. more than one order of magnitude larger) in the case of QL such as 4He, nuclei or electrons in atoms (see Table I). In the case of nuclei, 0.4, using typical values of the nucleon-nucleon interaction, namely V0 100 MeV and r0 1 fm [2]. The quantality indicates the importance to analyze the relation between the ZPE and the potential energy in finite systems. Moreover, among the quantality values of Table I, nuclei and atomic electrons have the largest one, indicating strong quantum effects. We shall discuss these 2 systems. It should be noted that i) quantality is valid for both fermions and bosons, ii) the quantum liquid nature of electron in atoms has been discussed in Ref. [2], and iii) results have to be considered within one order of magnitude, due to the semi-quantitative nature of the approach.
Constituent | m | r0 (nm) | V0 (eV) | T0 (eV) | State | ZPEv | QV | TI | ||||
20Ne atom | 20 | 0.31 | 31 10-4 | 1.1 10-5 | 4.9 10-6 | -6.1 1012 | 12.0 | 0.007 | crystal | 3.3 10-8 | 5.5 10-21 | 8.0 10-19 |
H2 molecule | 2 | 0.33 | 32 10-4 | 9.5 10-5 | 5.3 10-6 | -5.9 1011 | 4.1 | 0.06 | crystal | 3.2 10-7 | 5.3 10-19 | 9.0 10-18 |
4He atom | 4 | 0.29 | 8.6 10-4 | 6.0 10-5 | 1.2 10-6 | -4.4 1012 | 2.7 | 0.14 | QL | 1.9 10-7 | 4.7 10-20 | 2.7 10-19 |
3He atom | 3 | 0.29 | 8.6 10-4 | 0.8 10-5 | 1.2 10-6 | -3.2 1012 | 2.3 | 0.19 | QL | 2.6 10-7 | 7.7 10-20 | 3.8 10-19 |
Nucleon | 1 | 10-6 | 100 106 | 25 106 | 0.51 | 1.3 | 1.6 | 0.41 | QL | 0.18 | 0.02 | 0.04 |
e- in atoms | 5 10-4 | 0.05 | 10 | 16 | 2.5 10-3 | -5 104 | 0.6 | 3.1 | QL | 8.1 10-3 | 1.6 10-7 | 5.0 10-8 |
The inter-constituent distance r0 has a lower boundary rl, so that rrl : below this limit radius rl, the ZPE is large enough to overcome the potential energy and make the system unbound. The corresponding limit condition T0=V0 leads to
@@
Close to the limit case, such as for electrons in atoms, r0=rl and the quantality takes its maximal value, namely 2. This is confirmed by the phenomenological value of quantality [2], reproduced in Table I. In other systems, 2. For instance, in the case of nuclei, 1. This is a specific feature of nuclei: is of the order of unity, as well as their other dimensionless quantities. We define it as specific naturalness and its origin shall be discussed in section IV.1. For instance, although both atomic electrons and nuclei have a quantality value close to unity, it will not be the case for other dimensionless quantities in the case of atomic electrons, contrarily to nuclei, as discussed in section IV. Eq (1) also shows that monitors the delocalization effects with respect to the minimal radius rl: the larger the equilibrium distance r0, the smaller the quantality with respect to the maximal value 2.
Whether the particle is at equilibrium at a distance r0 equal to or larger than rl may depend on several causes, such as delocalization effects. In nuclei, using the typical magnitude V0 of the nucleon-nucleon interaction gives rl=0.5 fm from Eq. (3). This corresponds to the size of the repulsive core of the interaction. It can be understood with the specific naturalness of nuclear systems: as an order of magnitude reasoning, V0rc.
@@ III Quantality and the cluster phase in nuclei
In the case of nuclei, the spatial extension of the nucleonic wave function can be calculated microscopically from e.g. EDF approaches. It is well approximated by the confining length of a harmonic oscillator (HO) potential, which parameters are chosen to match the radius of a given nucleus [11]:
@@
Eq (11) shows that the typical nucleonic dispersion can be interpreted as the geometrical average of the limit radius from the ZPE and the nuclei one. It also emphasizes the contribution of both the quantality and finite-size effects on the dispersion. It should be noted that this equation can also be applied to atomic electrons, the HO confining potential being a first-order approximation to any bound system. Since both rl and R are close to the Bohr radius, as discussed in the previous section, the dispersion of the electron wave function is also close to the Bohr radius.
@@
Having identified the number of nucleons as a control parameter for the occurrence of a cluster phase, the role of the density in such phases can now be analyzed with the present approach. The inter-constituent distance is related to the density of the system. Let us consider a many-body system where the inter-constituent distance can change compared to its equilibrium value, due for instance, to external conditions such as compression or dilatation of the system, as it can occur in heavy-ion collisions. It can also be that the density of the system itself can be spatially dependent and hence have a lower density on the surface. For instance, finite nuclei display large fluctuations around , such that nucleons evolving in the low-density parts of finite nuclei could self-arrange into alpha-clusters [20, 21, 22, 23], as probed by recent (p,p) reactions [25].
@@
using Eq. (7). Hence, cluster states can occur in light nuclei (A10) at saturation density or at a lower density than the saturation one in heavy nuclei (A10): either the density or the number of nucleons can be a control parameter for the occurrence of cluster phase from a nuclear quantum liquid one. Interestingly, the threshold number of A 10 for clusterization is driven by the quantality. Eq. (16) also shows that in 120Sn, clusterization could occur below a density of about one-tenth of the saturation density, that is on its very surface. These results are in agreement with nuclear phenomenology, where lighter nuclei and/or smaller density (the so-called Mott transition [31, 32, 33, 34]) trigger cluster states. This is also in qualitative agreement with predictions of more microscopic approaches based e.g. on EDF, with or without explicit inclusion of alpha degrees of freedom [22, 21, 23], illustrating the relevance of the present discussion.
@@ IV.1 The key formula
where is the velocity of the constituents. Therefore, Eq (19) highlights the relation between the quantality, the coupling constant, and the velocity of the system.
@@
When there is no scalar mediator, such as in the vast majority of many-body systems, then =-1. A departure from this value indicates the presence of a scalar mediator in the interaction. For instance, in the case of nuclei, using the typical values of V and S leads to =1/90.1. is hereafter called the Pseudo Spin Symmetry (PSS) breaking coefficient because the =0 case corresponds to the realization of the PSS [36], namely V=-S. Moreover, the vector and the scalar potentials cannot be of the same sign, because of the repulsive/attractive nature of the spin triplet/singlet dependence of the interactions (check), implying 1. Therefore, the case =1 corresponds to the maximal PSS violating case.
@@
Let us discuss the above-mentioned dimensionless quantities on nuclei and atomic electrons. Nuclear states are known to usually behave as quantum liquid states and therefore 1, as displayed on Table I. Moreover, the low energy QCD interaction yields 1. Eq. (19) then provides 1. This value of allows for strong spin-orbit effect in finite systems [27]. The nucleus is therefore a very specific system where 1 within about one order of magnitude. This is due to naturalness.
In order to understand the origin of this specificity of naturalness in nuclei, the coupling constant Eq. (18) can be rewritten with the mass of the mediator m0cc/:
Therefore, naturalness could be interpreted by the close mass of the nucleon and the mesons, namely the 1 GeV scale, themselves related to the scale of chiral symmetry breaking and to the chiral condensate [37, 38]: in nuclei, the mass of the mediator is of the same order of magnitude than the constituent particle mass (mp/mπ 7 at most). The mm fact, responsible for 1, is related to QCD effects which provide a typical energy scale of few hundreds of MeV. This allows for similar masses between the lightest baryons and mesons. Eq. (26) exhibits the relationship between QCD driven masses ( 1), the strong spin-orbit effect in nuclei ( 1) and their QL behavior ( 1).
In the case of electrons in atoms, is close to 1 as well (even being close to its maximal value, see section II), which also behave as a QL [2]. However, this is not a specific naturalness system since the coupling constant is the electromagnetic one: =1/137, yielding 104 using Eq.(19). Namely, values are far from unity for and . Indeed, the definition of the spin-orbit parameter (Eq. (22)) yields mec2/V (105 eV/10 eV) 104, which is in agreement with the experimentally observed magnitude of the spin-orbit effect, compared to the shell one [27], i.e. the fine structure: 1/10-4.
@@ IV.3 Binding energy vs shell effects dependence on quantality
In both cases, approximated ways of solving the Schrödinger equation have been used (Fermi gas and HO).
@@ V Generalization of the quantality
The present study shows the pivotal role of the (18), (22) and (1) ratios to characterize several fundamental properties of many-body states. More generally, it is possible to build in a systematic way dimensionless quantities from the 3 basic quantities of the system: V0, r0 and m. New dimensionless quantities could be produced in the spirit of the quantality using the Pi theorem [43]. In the following, in order to have only energy units, and for the sake of simplicity, the r0 lengthscale can equivalently be described by the following mass:
@@
It corresponds to the velocity v0 of the constituents, due to the ZPE. Its values in various systems are given in Table I. In the case of nuclei, it has a non-negligible value compared to one. Eq (41) shows that it comes from the naturalness of this specific system. Following this interpretation, the velocity of the nucleons is non-negligible compared to the light speed (i.e. at the upper limit of the non-relativistic approximation) because the spin-orbit effect is important (1), as well as the magnitude of the interaction (1).
@@
It is, therefore, the ratio of the zero point motion velocity to the spin-orbit parameter. A large value corresponds to a large zero-point motion velocity and a large spin-orbit effect. This could be interpreted as a quadratic velocity quantity, namely the velocity of the system due to both zero-point motion and spin-orbit effect: as an approximation, let us consider the kinetic energy of the constituent of the order of V0, as it can be in liquids. In this case, ’2, where is the velocity of the constituent over c. Hence,
The typical values of the quadratic velocity for various systems are given in Table I. Once again, in nuclei, this value is closer to one and the largest.

### v2

# Use of quantality in nuclei and many-body systems

@@ I Introduction
Quantality, a powerful dimensionless ratio first introduced by de Boer [1], was used by Mottelson to indicate when a system behaves like a quantum liquid (QL) rather than a crystal. It is larger for quantum liquid states than for crystal ones [2, 3]. The former have constituents with delocalized wave-functions, leading to a homogeneous density, whereas the latter have localized constituents at given nodes. Superfluid Helium systems, nuclei, and atomic electrons are of QL type [2, 4, 5, 6]. More precisely, nuclei and atomic electrons are found to be the two most QL-like systems, having larger values of the quantality than superfluid Helium systems.
Let us consider a many-body system composed of constituents of mass m and interacting with an interaction of typical magnitude V0. The inter-constituent distance at equilibrium is r0. These 3 quantities characterize the main behavior of the system as they correspond to the input of its equation of motion. Instead of solving it as exactly as possible, a complementary way to predict the qualitative behavior of the system is to use indicators, incorporating step by step quantum mechanical effects. As a first step, the quantality is defined as the ratio of the Zero Point Energy (ZPE) T0 to the magnitude of the potential V0:
Quantality was used by Mottelson to characterize QL vs crystal behavior of systems [2, 3]. The typical value calculated with Eq. (1) is 10-2,-3 in the case of crystals like atoms and molecules, and 0.1-1 (i.e. more than one order of magnitude larger) in the case of QL such as 4He, nuclei or electrons in atoms (see Table I). In the case of nuclei, 0.4, using typical values of the nucleon-nucleon interaction, namely V0 100 MeV and r0 1 fm [2, 3]. The quantality indicates the importance to analyze the relation between the ZPE and the potential energy in finite systems. Moreover, among the quantality values of Table I, nuclei and atomic electrons have the largest one, indicating strong quantum effects. It should be noted that i) quantality is valid for both fermions and bosons, ii) the quantum liquid nature of electron in atoms has been discussed in Ref. [2], and iii) results have to be considered within one order of magnitude, due to the semi-quantitative nature of the approach.
| Constituent | m | r0 (nm) | V0 (eV) | T0 (eV) | State | ZPEv | QV | TI | ||||
| 20Ne atom | 20 | 0.31 | 31 10-4 | 1.1 10-5 | 0.007 | crystal | 4.9 10-6 | -6.1 1012 | 12.0 | 3.3 10-8 | 5.5 10-21 | 8.0 10-19 |
| H2 molecule | 2 | 0.33 | 32 10-4 | 9.5 10-5 | 0.06 | crystal | 5.3 10-6 | -5.9 1011 | 4.1 | 3.2 10-7 | 5.3 10-19 | 9.0 10-18 |
| 4He atom | 4 | 0.29 | 8.6 10-4 | 6.0 10-5 | 0.14 | QL | 1.2 10-6 | -4.4 1012 | 2.7 | 1.9 10-7 | 4.7 10-20 | 2.7 10-19 |
| 3He atom | 3 | 0.29 | 8.6 10-4 | 0.8 10-5 | 0.19 | QL | 1.2 10-6 | -3.2 1012 | 2.3 | 2.6 10-7 | 7.7 10-20 | 3.8 10-19 |
| Nucleon | 1 | 10-6 | 100 106 | 25 106 | 0.41 | QL | 0.51 | 1.3 | 1.6 | 0.18 | 0.02 | 0.04 |
| e- in atoms | 5 10-4 | 0.05 | 10 | 16 | 3.1 | QL | 2.5 10-3 | -5 104 | 0.6 | 8.1 10-3 | 1.6 10-7 | 5.0 10-8 |
@@
Despite its powerfulness, the role of quantality has been very scarcely studied in nuclear physics [2, 3, 5, 16, 6]. In this work, we further explore how quantality can be used to understand the main differences between atomic electrons and nucleons in their quantum liquid phases. It is then used to identify the control parameters for the cluster phase in nuclei, namely the role of the number of nucleons and the density. The relation to the low-density transition towards clusterization is analyzed. The use of quantality, as a dimensionless ratio, to understand its relation with the interaction coupling constants leads to a fundamental relation, allowing to study the magnitude of the spin-orbit effect in finite nuclei. The relative magnitude of the binding energy to the shell-structure one is also shown to depend on the quantality.
The present work aims to analyze the behavior of many-body systems with simple but powerful indicators, by progressively refining their quantum description. First, the role of the ZPE is discussed. Then, the inclusion of the dispersion of the wave-function of the constituents, from the solution of a e.g. Schrödinger equation is considered. As a third step, the relation with the spin-orbit effect is given, based on the non-relativistic reduction of the Dirac equation. It will be shown how quantality plays a pivotal role during all these three stages. In section II, the emergence of a fundamental lengthscale, the limit radius, is studied in relation to the ZPE and the quantality. The relation between the quantality and the localization parameter is discussed in section III, showing how the appearance of clusters in the low-density regime is recovered. Section IV relates the quantality to the dimensionless magnitude of the interaction, showing how the spin-orbit effect in finite systems can be deduced. Finally, a systematic study of possible dimensionless quantities is provided in Section V, generalizing the concept of quantality.
@@ II The quantality, the zero point energy, and the limit radius
It should be noted that in a more general manner, the quantality could be defined as the ratio of the mean values of the kinetic energy to the potential one:
where Eq. (1) is the application to the phenomenological case with a given potential and kinetic energy value. Eq. (3) allows the calculation of the value of the quantality for any interaction and many-body method, by using the corresponding wave functions and operators. Such a control parameter might prove useful in Ab Initio approaches. For instace, it could be possible to monitor the degree of clusterisation as a function of the similarity renormalisation group evolution of the bare chiral potential [30, 31, 32], or of the many-body scheme at stake [33, 34, 20] . One can then study with these approaches if, e.g., the quantum liquid behavior of the system is in agreement with the calculated value. In the present work, we will keep the use of Eq. (1), as it already paves the way for numerous developments.
Starting with the inter-constituent distance r0, it has a lower boundary rl, so that rrl : below this limit radius rl, the ZPE is large enough to overcome the potential energy and make the system unbound. The corresponding limit condition T0=V0 leads to
@@
Close to the limit case, such as for electrons in atoms, r0=rl and the quantality takes its maximal value, namely 2. This is confirmed by the phenomenological value of quantality [2, 3], reproduced in Table I. In other systems, 2. For instance, in the case of nuclei, 1. This is a specific feature of nuclei: is of the order of unity, as well as their other dimensionless quantities. We define it as specificness and its origin shall be discussed in section IV.1. For instance, although both atomic electrons and nuclei have a quantality value close to unity, it will not be the case for other dimensionless quantities in the case of atomic electrons, contrarily to nuclei, as discussed in section IV. Eq (1) also shows that monitors the delocalization effects with respect to the minimal radius rl: the larger the equilibrium distance r0, the smaller the quantality with respect to the maximal value 2.
Whether the particle is at equilibrium at a distance r0 equal to or larger than rl may depend on several causes, such as delocalization effects. In nuclei, using the typical magnitude V0 of the nucleon-nucleon interaction gives rl=0.5 fm from Eq. (4). This corresponds to the size of the repulsive core of the interaction. It can be understood with the specificness of nuclear systems: as an order of magnitude reasoning, V0rc. Put in other words, the rl=0.5 fm short-scale distance corresponds to an energy of c/ 400 MeV which defines an energy cutoff above which the ZPE is too large for the system to be bound. Interestingly, this about 400 MeV value also corresponds to the phenomenological determination of the cutoff parameter in chiral EFT [35, 36], pointing to a possible link with the ZPE limit, which remains to be investigated.
@@ III Quantality and the cluster phase in nuclei
In the case of nuclei, the spatial extension of the nucleonic wave function can be calculated microscopically from e.g. EDF approaches [14]. It can also be calculated from ab-initio or lattice EFT approaches for instance, by computing the spatial dispersion r of the corresponding wave functions. Similarly, the interconstituent distance could be computed from lattice-EFT calculations, in order to obtain a microscopic value. Here, in order to analyse the main impacts of the localization parameter, we use the sound approximation by the confining length of a harmonic oscillator (HO) potential, which parameters are chosen to match the radius of a given nucleus [14]:
@@
Eq (12) shows that the typical nucleonic dispersion can be interpreted as the geometrical average of the limit radius from the ZPE and the nuclei radius. It also emphasizes the contribution of both the quantality and finite-size effects on the dispersion. It should be noted that this equation can also be applied to atomic electrons, the HO confining potential being a first-order approximation to any bound system. Since both rl and R are close to the Bohr radius, as discussed in the previous section, the dispersion of the electron wave function is also close to the Bohr radius.
@@
Having identified the number of nucleons as a control parameter for the occurrence of a cluster phase, the role of the density in such phases can now be analyzed with the present approach. The inter-constituent distance is related to the density of the system. Let us consider a many-body system where the inter-constituent distance can change compared to its equilibrium value, due for instance, to external conditions such as compression or dilatation of the system, as it can occur in heavy-ion collisions. It can also be that the density of the system itself is spatially dependent and hence have a lower density on the surface. For instance, finite nuclei display large fluctuations around , such that nucleons evolving in the low-density parts of finite nuclei could self-arrange into alpha-clusters [23, 24, 25, 26], as probed by recent (p,p) reactions [28].
@@
using Eq. (8). Hence, cluster states can occur in light nuclei (A10) at saturation density or at a lower density than the saturation one in heavy nuclei (A10): either the density or the number of nucleons can be a control parameter for the occurrence of cluster phase from a nuclear quantum liquid one. Interestingly, the threshold number of A 10 for clusterization is driven by the quantality. Eq. (17) also shows that in 120Sn, clusterization could occur below a density of about one-tenth of the saturation density, that is on its very surface. These results are in agreement with nuclear phenomenology, where lighter nuclei and/or smaller density (the so-called Mott transition [40, 41, 42, 43]) trigger cluster states. This is also in qualitative agreement with predictions of more microscopic approaches based e.g. on EDF, with or without explicit inclusion of alpha degrees of freedom [25, 24, 26], illustrating the relevance of the present discussion. Fig. 2 summarizes the conditions for the cluster and quantum liquid phases induced by the condition (17), exhibiting both the density and the number of nucleons as control parameters. This picture, validated in the relativistic EDF, could be further investigated in other approaches such as ab-initio and EFT ones.
@@
It should be noted that both the small value of the quantality directly obtained from the alpha-alpha potential [4] and the present analysis of the dimensionless ratio of the typical alpha-particle size to the internucleon distance, point to alphas as possible degrees of freedom of a dedicated EFT.
@@ IV.1 The key formula
where is the velocity of the constituents. Therefore, Eq (20) highlights the relation between the quantality, the coupling constant, and the velocity of the system. For instance, it could be used, in an ab-initio or EFT approach, to calculate the quantality from the behavior of the coupling constant and using V instead of V0 as in Eq. (3):
@@
When there is no scalar mediator, such as in the vast majority of many-body systems, then =-1. A departure from this value indicates the presence of a scalar mediator in the interaction. For instance, in the case of nuclei, using the typical values of V and S leads to =1/90.1. is hereafter called the Pseudo Spin Symmetry (PSS) breaking coefficient because the =0 case corresponds to the realization of the PSS [45], namely V=-S. Moreover, the vector and the scalar potentials cannot be of the same sign, because of the repulsive/attractive nature of the spin triplet/singlet dependence of the interactions, implying 1. Therefore, the case =1 corresponds to the maximal PSS violating case.
@@
Let us discuss the above-mentioned dimensionless quantities on nuclei and atomic electrons. Nuclear states are known to usually behave as quantum liquid states and therefore 1, as displayed in Table I. Moreover, the low energy QCD interaction yields 1. Eq. (20) then provides 1. This value of allows for strong spin-orbit effect in finite systems [7]. The nucleus is therefore a very specific system where 1 within about one order of magnitude, as seen in Table I. We define this property as specificness. It should be noted that this definition differs from the naturalness one used in EFT, where renormalized dimensionless low energy constants are [46].
In order to understand the origin of this specificness in nuclei, the coupling constant Eq. (19) can be rewritten with the mass of the mediator m0cc/:
Therefore, specificness could be interpreted by the close mass of the nucleon and the mesons, namely the 1 GeV scale, themselves related to the scale of chiral symmetry breaking and to the chiral condensate [47, 48]: in nuclei, the mass of the mediator is of the same order of magnitude than the constituent particle mass (mp/mπ 7 at most). The mm fact, responsible for 1, is related to QCD effects which provide a typical energy scale of few hundreds of MeV. This allows for similar masses between the lightest baryons and mesons. Eq. (28) exhibits the relationship between QCD driven masses ( 1), the strong spin-orbit effect in nuclei ( 1) and their QL behavior ( 1). In other words, nuclei are specific in the sense that mainly one interaction, namely the strong force, is involved in both their constitution and their interactions.
In the case of electrons in atoms, is close to 1 as well (even being close to its maximal value, see section II), which also behave as a QL [2]. However, this is not a specificness system since the coupling constant is the electromagnetic one: =1/137, yielding 104 using Eq.(20). Namely, values are far from unity for and . Indeed, the definition of the spin-orbit parameter (Eq. (24)) yields mec2/V (105 eV/10 eV) 104, which is in agreement with the experimentally observed magnitude of the spin-orbit effect, compared to the shell one [7], i.e. the fine structure: 1/10-4.
@@ IV.3 Binding energy vs shell effects dependence on quantality
It should be noted that the Fermi gas and the HO both correspond to approximate ways of solving the Schrödinger equation.
@@ V Generalization of the quantality
The present study shows the pivotal role of the (19), (24) and (1) dimensionless ratios to characterize several fundamental properties of many-body states. More generally, it is possible to build in a systematic way dimensionless quantities from the 3 basic quantities of the system: V0, r0 and m. New dimensionless quantities could be produced in the spirit of the quantality using the Pi theorem [53]. It should be noted, in a general way, that the value of dimensionless parameters allows the identification of the regime where a given phenomenon (clusterization, quantum liquid behavior, etc.) is dominant and to build an EFT that describes this regime. In the following, in order to have only energy units, and for the sake of simplicity, the r0 lengthscale can equivalently be described by the following mass:
@@
It corresponds to the velocity v0 of the constituents, due to the ZPE. Its values in various systems are given in Table I. In the case of nuclei, it has a non-negligible value compared to one. Eq (43) shows that it comes from the specificness of this system. Following this interpretation, the velocity of the nucleons is non-negligible compared to the light speed (i.e. at the upper limit of the non-relativistic approximation) because the spin-orbit effect is important (1), as well as the magnitude of the interaction (1).
@@
Therefore, QV is the ratio of the zero point motion velocity to the spin-orbit parameter. A large value corresponds to a large zero-point motion velocity and a large spin-orbit effect. In addition to the ZPE velocity, QV can be also related to the quadratic velocity of the constituents of the system: as an approximation, let us consider the kinetic energy of the constituent of the order of V0, as it can be in liquids. In this case, (22) gives ’2/, where is the velocity of the constituent over c. Hence,
The typical values of the quadratic velocity for various systems are given in Table I. Once again, in nuclei, this value is the closest to one and the largest, compared to other systems.